OpenStudy (anonymous):

Solve using the elimination method 6x - 10y = -18 -3x + 2y = 3

OpenStudy (anonymous):

@BunnyBree

OpenStudy (anonymous):

The first thing we need to do is eliminate one of the variables. I'm going to eliminate x this time. So I have to multiply the bottom equation by 2. 6x - 10y = -18 (-3x + 2y = 3) x 2 We get: 6x - 10y = -18 -6x + 4y = 6 Add the two equations together. -6y = -12 Divide each side by -6. y = 2 <--- One of the solutions. Now plug y into one of the original equations to obtain x.

OpenStudy (anonymous):

I'm going to use the first equation. 6x - 10y = -18 6x - 10(2) = -18 6x - 20 = -18 Add 20 to both sides. 6x = 2 Divide each side by 6. x = 1/3 The solutions are: x = 1/3, and y = 2.